# MMedAgent-RL: Optimizing Multi-Agent Collaboration for Multimodal Medical Reasoning

### v1

@@ Abstract
Medical Large Vision-Language Models (Med-LVLMs) have shown strong potential in multimodal diagnostic tasks. However, existing single-agent models struggle to generalize across diverse medical specialties, limiting their performance. Recent efforts introduce multi-agent collaboration frameworks inspired by clinical workflows, where general practitioners (GPs) and specialists interact in a fixed sequence. Despite improvements, these static pipelines lack flexibility and adaptability in reasoning. To address this, we propose MMedAgent-RL, a reinforcement learning (RL)-based multi-agent framework that enables dynamic, optimized collaboration among medical agents. Specifically, we train two GP agents based on Qwen2.5-VL via RL: the triage doctor learns to assign patients to appropriate specialties, while the attending physician integrates the judgments from multi-specialists and its own knowledge to make final decisions. To address the inconsistency in specialist outputs, we introduce a curriculum learning (CL)-guided RL strategy that progressively teaches the attending physician to balance between imitating specialists and correcting their mistakes. Experiments on five medical VQA benchmarks demonstrate that MMedAgent-RL not only outperforms both open-source and proprietary Med-LVLMs, but also exhibits human-like reasoning patterns. Notably, it achieves an average performance gain of 18.4% over supervised fine-tuning baselines.
@@ 1 Introduction
The primary contribution of this paper is MMedAgent-RL, an RL-driven framework optimized for multi-agent collaboration in improving medical reasoning. Empirical results on five medical multimodal datasets, shows that the model performs exceptionally well not only on in-domain datasets but also on out-of-domain datasets, outperforming a series of both open-source and proprietary LVLMs, exceeding SFT method [40] by 18.4%, and our 3B model rivals the performance of a 7B model. In addition to the quantitative performance, we also observed that the model’s reasoning process resembles human doctors’ thought processes: it first defines the disease mentioned in the question, carefully analyzes the image, and then determines if it can fully match the disease definition.
@@ 4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL [3] as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Appendix. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
@@ 4.2 Main Results
Performance in Out-of-Distribution Datasets. We evaluate the performance of MMedAgent-RL across various out-of-distribution (OOD) datasets. The results are presented in Table 1, which demonstrates the generalization of our approach in adapting to different OOD scenarios. These two OOD datasets cover multiple body parts and involve various medical image modalities. Through reinforcement learning, MMedAgent-RL demonstrates significant superiority across multiple modalities, outperforming the base model by 20% and the SFT method by 20.2%. Moreover, it surpasses the performance of multi-agent collaboration methods that cannot optimize models, i.e., MedAgents and MDAgents, by 13% and 8%, highlighting the effectiveness of our approach in handling diverse and unseen data distributions.
@@ 4.3 Analysis
Case Study and Outlook on “Aha Moments". As shown in Figure 6, we demonstrate the model’s performance on multiple cases. We can clearly observe that MMedAgent-RL can provide accurate answers in the <answer> tags and also generate high-quality reasoning processes. We also found that this resembles the thought process of human doctors: it first defines the disease mentioned in the problem, carefully analyzes the images, and then determines whether it completely matches the disease definition. Moreover, it also analyzes the results provided by specialists. In our cases, it can accurately evaluate the specialists’ outputs and then reason out the correct answer. Although there isn’t an “aha moment” seen in LLMs [13], this structured reasoning may inspire more human-like scientific AI systems in the future.
@@ E.4 More Cases
To further demonstrate the robustness and versatility of our proposed model in multimodal medical applications, we present additional representative cases in Figure 8, Figure 9, Figure 10 and Figure 11. These examples encompass various clinical scenarios and imaging modalities, providing a comprehensive illustration of the model’s ability to effectively integrate and interpret diverse types of medical data.
Figure 8: Case 1 (CT scan of foot): The model identifies plantar fascia pathology from a CT image, requiring anatomical knowledge of soft tissue structures in the foot and recognition of characteristic changes consistent with plantar fasciitis. Case 2 (Protein structure diagram): A structural biology reasoning task based on a protein diagram, where the absence of distinct subunit boundaries leads to the conclusion that the protein does not exhibit quaternary structure—showcasing visual-structural reasoning in molecular biology.
Figure 9: Case 3 (Ultrasound image): The model interprets an ultrasound to identify the pancreas, integrating grayscale image features with organ shape and texture—an example of soft tissue recognition using sonography. Case 4 (Brain MRI with question about vestibular system): This task highlights the model’s ability to identify mismatches between the image modality (brain MRI) and the clinical question (vestibular system), and to reason that facial sensation loss is not explainable from the given scan.
Figure 10: Case 5 (Abdominal CT image – polysplenia diagnosis): The task involves recognizing abnormal organ positioning indicative of polysplenia. This case highlights anatomical reasoning through CT imaging, requiring recognition of fused or duplicated spleens and an understanding of associated syndromic features. Case 6 (Histopathology slide – adenoma diagnosis): A histological image is used to assess the presence of malignant features. The model correctly distinguishes normal plantar fascia morphology from pathological adenoma patterns, demonstrating reasoning in pathology image interpretation.
Figure 11: Case 7 (Pie chart comparison – elemental composition): This case blends image data (a pie chart of Earth’s element composition) with textual biochemical data (composition of the human body). The reasoning required crosses domains, comparing datasets to identify overlapping elements, exemplifying multimodal cross-referencing and synthesis.

### v2

@@ Abstract
Medical Large Vision-Language Models (Med-LVLMs) have shown strong potential in multimodal diagnostic tasks. However, existing single-agent models struggle to generalize across diverse medical specialties, limiting their performance. Recent efforts introduce multi-agent collaboration frameworks inspired by clinical workflows, where general practitioners (GPs) and specialists interact in a fixed sequence. Despite improvements, these static pipelines lack flexibility and adaptability in reasoning. To address this, we propose MMedAgent-RL, a reinforcement learning (RL)-based multi-agent framework that enables dynamic, optimized collaboration among medical agents. Specifically, we train two GP agents based on Qwen2.5-VL via RL: the triage doctor learns to assign patients to appropriate specialties, while the attending physician integrates the judgments from multi-specialists and its own knowledge to make final decisions. To address the inconsistency in specialist outputs, we introduce a curriculum learning (CL)-guided RL strategy that progressively teaches the attending physician to balance between imitating specialists and correcting their mistakes. Experiments on five medical VQA benchmarks demonstrate that MMedAgent-RL not only outperforms both open-source and proprietary Med-LVLMs, but also exhibits human-like reasoning patterns. Notably, it achieves an average performance gain of 20.7% over supervised fine-tuning baselines.
@@ 1 Introduction
The primary contribution of this paper is MMedAgent-RL, an RL-driven framework optimized for multi-agent collaboration in improving medical reasoning. Empirical results on five medical multimodal datasets, shows that the model performs exceptionally well not only on in-domain datasets but also on out-of-domain datasets, outperforming a series of both open-source and proprietary LVLMs, exceeding SFT method [40] by 20.7%. In addition to the quantitative performance, we also observed that the model’s reasoning process resembles human doctors’ thought processes: it first defines the disease mentioned in the question, carefully analyzes the image, and then determines if it can fully match the disease definition.
@@ 4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL [3] (3B and 7B) as the base model. We design the prompt template using the format employed in MM-EUREKA [35], clearly specifying the required output structure, which includes using <think> and <answer> tags to separately contain the reasoning process and the final answer, with the two being separated. The detailed prompt is shown in Appendix. For training hyperparameters, the rollout batch size and training batch size are both set to 128, with 8 rollouts generated for each sample. The sampling temperature is set to 1.0 to encourage response diversity, and optimization is done with a learning rate of $1\times 10^{-6}$. Additionally, for the three stages of curriculum reinforcement learning, the KL divergence coefficients are set to $1\times 10^{-3}$, $4\times 10^{-3}$, and $1\times 10^{-2}$ respectively to stabilize training. For the number of specialists, we set $e=3$. For the framework, we adopt a multimodal RL framework based on OpenRLHF [17]. All training is conducted on 8 NVIDIA Tesla A100 80GB GPUs.
@@ 4.2 Main Results
Performance in Out-of-Distribution Datasets. We evaluate the performance of MMedAgent-RL across various out-of-distribution (OOD) datasets. The results are presented in Table 1, which demonstrates the generalization of our approach in adapting to different OOD scenarios. These two OOD datasets cover multiple body parts and involve various medical image modalities. Through reinforcement learning, MMedAgent-RL demonstrates significant superiority across multiple modalities, outperforming the base model by 15.3% and the SFT method by 25.1%. Moreover, it surpasses the performance of multi-agent collaboration methods that cannot optimize models, i.e., MedAgents and MDAgents, by 13% and 8%, highlighting the effectiveness of our approach in handling diverse and unseen data distributions.
@@ 4.3 Analysis
Case Study and Outlook on “Aha Moments". As shown in Figure 6, we demonstrate the model’s performance on multiple cases. We can clearly observe that MMedAgent-RL can provide accurate answers in the <answer> tags and also generate high-quality reasoning processes. We also found that this resembles the thought process of human doctors: it first defines the disease mentioned in the problem, carefully analyzes the images, and then determines whether it completely matches the disease definition. Moreover, it also analyzes the results provided by specialists. In our cases, it can accurately evaluate the specialists’ outputs and then reason out the correct answer. Please refer to more cases in Sec. E.4, with a greater variety of examples. Although there isn’t an “aha moment” seen in LLMs [13], this structured reasoning may inspire more human-like scientific AI systems in the future.
@@ E.4 More Cases
To further demonstrate the robustness and versatility of our proposed model in multimodal medical applications, we present additional representative cases in Figure 8, Figure 9, Figure 10, Figure 11, Figure 12 and Figure 13. These examples encompass various clinical scenarios and imaging modalities, providing a comprehensive illustration of the model’s ability to effectively integrate and interpret diverse types of medical data.
Figure 8: Case 1 (OmniMedVQA) & Case2 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 9: Case 3 (OmniMedVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 10: Case 4 (PathVQA). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 11: Case 5 (VQA-RAD) & Case 6 (VQA-RAD). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 12: Case 7 (PathVQA) & Case 8 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.
Figure 13: Case 9 (MMMU). In the model responses, blue text represents the process of reasoning about relevant medical knowledge based on the question, and red text represents the analysis of the answer provided by the specialists.